what is the width of the rectangle if the length is 3 times the width and the premeter is 240
Answer by mando271(42)

to find the perimeter of a rectangle you need 2 lenghts and 2 widths, so the formula is:
2L + 2W = P
in our problem, we are trying to find the width so you substitute the givens:
2(3W) + 2W = 240
now you can solve for the width
6W + 2W = 240
8W = 240
W = 30